OpenStudy (anonymous):

What is the Mean of 2,6,15,3,1,8,7

OpenStudy (anonymous):

add them all up and divide by the how many numbers there are, this set has n=7 numbers so you take the sum of all elements and divide by n

OpenStudy (anonymous):

Thank You :)

OpenStudy (anonymous):

42/7=6

OpenStudy (anonymous):

K

OpenStudy (anonymous):

34+57+68+12+9=

OpenStudy (anonymous):

what

OpenStudy (anonymous):

34+57+68+12+9=

OpenStudy (anonymous):

what does that have to do with your question?

OpenStudy (anonymous):

Ya and the answer id 36 right?

OpenStudy (anonymous):

how can you add stuff to 57 and get a number less than 57 i think you subtracted in there

OpenStudy (anonymous):

what? i'm confused

OpenStudy (anonymous):

its 180 btw

OpenStudy (anonymous):

do you know how to add?

OpenStudy (anonymous):

YA 7th grader

OpenStudy (anonymous):

ok when you have this, 34+57+68+12+9 explain to me how u got 36

OpenStudy (anonymous):

I not know how i got it

OpenStudy (anonymous):

well you must know because you did something to get 36 unless you just guessed

OpenStudy (anonymous):

Is it 405.4

OpenStudy (anonymous):

do you think this is a joke?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

people go on here to get actual help

OpenStudy (anonymous):

I know and i need help

OpenStudy (anonymous):

put this in a calculator 34+57+68+12+9

OpenStudy (anonymous):

180

OpenStudy (anonymous):

and see what happens thats all i can say, if you have 68 apples and add apples to them you can never gt less than 68 apples

OpenStudy (anonymous):

i still get 36

OpenStudy (anonymous):

then its hopeless to try and help you im sorry

OpenStudy (anonymous):

What did u get

OpenStudy (anonymous):

?

OpenStudy (anonymous):

180...

OpenStudy (anonymous):

then Divide by 5

OpenStudy (anonymous):

well you only said 34+57+68+12+9 which equals 180 you never said to find the mean of those numbers lol

OpenStudy (anonymous):

but yes you would divide by 5 and get 36.....

OpenStudy (anonymous):

But do not u have to dived the numbers

OpenStudy (anonymous):

mean is average yes you have to divide. if you have 2 tests in a course and u have 50 and 80 on them you add 50+80=130 and divide by 2 to find your average

OpenStudy (anonymous):

same with this stuff

OpenStudy (anonymous):

The answer is 36

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

haha i wans solving that yesterday like that kind of problems